OpenStudy (anonymous):

Simplify the following expression: 5 - [(3i + 4) - 2i]

jagr2713 (jagr2713):

first we Simplify 5-(3 i+4-2 i) Distribute -1 over 4+i. -(4+i) = -4-i: -i-4+5 Evaluate -4-i+5. -4-i+5 = (5-4)-i = 1-i so we would get 1-i

OpenStudy (anonymous):

OMG thank you so much!

jagr2713 (jagr2713):

Yw :D